Need help on this problem...
16x^-4 - 65x^-2 +4 = 0
where u = x^-2
So far I've done...
16u^2 - 65u + 4 = 0
(16u - 1)(u - 4) = 0
u = 1/16 and u = 4
From there, I need help.

Answer by rothauserc(4718)

u = 1/16 and u = 4
picking up where you left off, note that u = x^-2
1) 1/x^2 = 1/16
x^2 = 16
x = 4 or-4
2) 1/x^2 = 4/1
4x^2 = 1
x^2 = 1/4
x = 1/2 or -1/2
now check the answers
16x^-4 - 65x^-2 +4 = 0
16/x^4 - 65/x^2 +4 = 0
now try x = 4
16/256 - 65/16 +4 = 0
0.0625 - 4.0625 +4 = 0
-4 +4 = 0
0 = 0
now try x = 1/2
16/(1/16) -65/(1/4) +4 = 0
-4 +4 = 0
0 = 0
the answers check